OpenStudy (anonymous):

imaginary number pattern equation? How can I simplify this without expanding it the long way...? (i^68)

OpenStudy (anonymous):

what is i^2?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

use rules of exponents to change i^68 to (i^2)^something

OpenStudy (anonymous):

but how would that work

OpenStudy (anonymous):

like I know youcan do (i^17)^4 but how do you know the answer is 1

jimthompson5910 (jim_thompson5910):

One shortcut is to divide 68 by 4 to get a quotient and a remainder if the remainder is 0, then i^68 = 1 if the remainder is 1, then i^68 = i if the remainder is 2, then i^68 = -1 if the remainder is 3, then i^68 = -i

OpenStudy (anonymous):

ah. change it around to (i^4)^17 what is i^4? what is (-1)^2?

OpenStudy (anonymous):

OH THAT IS SO COOL LOL @jim_thompson5910 THNX

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 wait but what if its not divisible by 4

jimthompson5910 (jim_thompson5910):

that's why we only care about the remainder

OpenStudy (anonymous):

OH ahha.. I see thanks

jimthompson5910 (jim_thompson5910):

for example, i^9 = i because 9/4 = 2 remainder 1 i^15 = -i since 15/4 = 3 remainder 3 i^60 = 1 since 60/4 = 15 remainder 0 we're only looking at the remainders when we're using that rule

OpenStudy (anonymous):

alright I understand it thanks!

jimthompson5910 (jim_thompson5910):

np